A particle is moving in x-y plane. It's initial velocity & acceleration are u=(4î+5ĵ) m/s and a=(2î-4ĵ) m/s^2. Find
(1) the time when the particle will cross the x axis
(2) x coordinate of the particle at that instant
(3) velocity of the particle at this instant
Initial coordinates of particle are (4m,10m)
1
Expert's answer
2019-09-23T09:19:19-0400

Question (a)

The cutting of the X axis occurs when the Y coordinate is equal to zero.


The vertical position is given by:"Y=Y_{o}+V_{oy}*t+\frac{1}{2}a_{y}t^{2}"


Where:

  • Initial position. "Y_{o}=10m"
  • Initial velocity "V_{oy}=+5\frac{m}{s}"
  • Vertical acceleration. a"a_{y}=-4\frac{m}{s^{2}}"

Evaluating at Y = 0.


"0=10+5t-\frac{1}{2}4t^{2}\\ -2t^{2}+5t+10=0"


A second grade equation."-2t^{2}+5t+10=0->ax^{2}+bx+c=0"


Coefficients "a=-2;b=5;c=10"


Solving the second degree equation.


"t=\frac{-b\pm \sqrt{b^{2}-4*a*c}}{2*a} \\t=\frac{-(5)\pm \sqrt{(5)^{2}-4*-2*10}}{2*-2}\\ t=\frac{-5 \pm 10.2}{-4}"


Solution 1

"t=\frac{-5 + 10.2}{-4} \\t=-1.25s"


Solution 2

"t=\frac{-5 - 10}{-4} \\t=3.75s"


the time when the particle will cross the x axis 

"\boxed{t=3.75s}"


Question (b)


The position of the particle is given by:


"\vec{r(t)}=\vec{r}_{o}+\vec{V_{o}}*t+\frac{1}{2}\vec{a}t^{2}"


Where:

  • initial velocity "\vec{V_{0}}=(4i+5j)\frac{m}{s}"
  • The acceleration is "\vec{a}=(2i-4j)\frac{m}{s^{2}}"
  • Initial position "\vec{r_{0}}=(4i+10j)m"
  • time "t=3.75s"


Numerically evaluating


"\vec{r(t)}=(4i+10j)m+(4i+5j)\frac{m}{s}*3.75s+\frac{1}{2}*(2i-4j)\frac{m}{s^{2}}*(3.75s)^{2}"

"\vec{r_{o}}=(4i+10j)m+(15i+18j)m+(14i-28j)m"

"\vec{r_{o}}=(33i+0j)m"


The position at that moment is:

"\boxed{\vec{r_{o}}=(33i+0j)m \text{ or } \vec{r_{o}}=(33m,0m)}"


Question (c)


The velocity of the particle is given by:


"\vec{V(t)}=\vec{V_{o}}+\vec{a}*t"


Where:

  • The initial velocity"\vec{V_{0}}=(4i+5j)\frac{m}{s}"
  • Acceleration "\vec{a}=(2i-4j)\frac{m}{s^{2}}"
  • Time t=3.75s


Numerically evaluating


"\vec{V(t)}=(4i+5j)\frac{m}{s} +(2i-4j)\frac{m}{s^{2}}*3.75s \\\vec{V(t)}=(4i+5j)\frac{m}{s}+(5.5i-15j)\frac{m}{s}\\\vec{V(t)}=(11.5i-10j)\frac{m}{s}"


The velocity at that moment is

"\boxed{vec{V(t)}=(11.5i-10j)\frac{m}{s} }"